40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12104     Entry time: Mon May 2 19:14:18 2016     In reply to: 12099     Reply to this: 12105   12122
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: Optical layout almost complete 

With Steve's help, I installed the Oplev earlier today. I adjusted the positions of the two lenses until I deemed the spot size on the QPD satisfactory by eye. As a quick check, I verified using the DTT template that the UGF is ~5Hz for both pitch and yaw. There is ~300uW of power incident on the QPD (out of ~2mW from the HeNe). In terms of ADC counts, this is ~13,000 counts which is about what we had prior to taking the endtable apart. There are a couple of spots from reflections off the black glass plate in the vacuum chamber, but in general, I think the overall setup is acceptable.

This completes the bulk of the optical layout. The only bits remaining are to couple the IR into the fiber and to install a power monitoring PD. Pictures to follow shortly. 

Now that the layout is complete, it remains to optimize various things. My immediate plan is to do the following:

  1. Maximize green transmission by tweaking alignment. I should also do a quick check using mirror specs to see that the measured transmitted green power compares favourably to what is expected.
  2. Check the green PDH loop transfer function at the X end - this will allow me to set the gain on the uPDH box systematically.
  3. Re-establish green beats, check noise performance.
  4. There are possibly multiple beam dumps that have to be installed. For now, I've made sure that no high power IR beams are incident on the enclosure. But there are a couple of red and green beams that have to be accounted for.

I will also need to upload the layout drawing to reflect the layout finally implemented.


Not directly related:

The ETMx oplev servo is now on. I then wanted to see if I could lock both arms to IR. I've managed to do this successfully - BUT I think there is something wrong with the X arm dither alignment servo. By manually tweaking the alignment sliders on the IFOalign MEDM screen, I can get the IR transmission up to ~0.95. But when I run the dither, it drives the transmission back down to ~0.6, where it plateaus. I will need to investigate further. 

 

GV Edit: There was some confusion while aligning the Oplev input beam as to how the wedge of the ETM is oriented. We believe the wedge is horizontal, but its orientation (i.e. thicker side on the right or left?) was still ambiguous. I've made a roughly-to-scale sketch (attachment #1) of what I think is the correct orientation - which turns out to be in the opposite sense of the schematic pinned up in the office area.. Does this make sense? Is there some schematic/drawing where the wedge orientation is explicitly indicated? My search of the elog/wiki did not yield any..

Attachment 1: ETMX_wedge.pdf  28 kB  Uploaded Mon May 2 23:57:49 2016  | Hide | Hide all
ETMX_wedge.pdf
ELOG V3.1.3-